Select two variables, set up two equations and solve the problem below. The sum of two integers is 45. The larger is 3 less than three times the smaller. Find the two integers.
How do I find and work this problem please give details each step so I can work and understand how it's done. Thank you

Answer by lwsshak3(11628)

Select two variables, set up two equations and solve the problem below. The sum of two integers is 45. The larger is 3 less than three times the smaller. Find the two integers.
***
let x=larger integer
let y=smaller integer
..
x+y=45 (sum of two integers is 45)
x=45-y
x=3y-3 (larger is 3 less than three times the smaller)
45-y=3y-3
4y=48
y=12
x=45-y=33
larger integer=33
smaller integer=12